OpenStudy (anonymous):

he is eligible for health insurance after he works 1,500 hours . last week had worked 1,441.75 hours , this week he worked 23.5hours . how many more hours must he work to be qualify ?

OpenStudy (anonymous):

He needs to work for more 34.75 hours.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

1441.75 + 23.5 + x = 1500( Where x is the number of hours he still needs to work) x = 1500 - 1465.5 Hence x = 34.75 hours.

OpenStudy (anonymous):

You're welcome.